# Sub-keV dark matter can strongly ionize molecular clouds

### v1

@@ Abstract
We show that the ionization of dense molecular clouds can be used to set strong constraints on dark matter models producing UV/X-ray photons in their annihilation or decay. We place robust and competitive constraints on various dark matter models, such as axion-like particles, scalars and sterile neutrinos, for masses between eV and keV, and project forecasts to illustrate the potential of this target. We discuss how these constraints can be significantly improved by considering a more refined sample of molecular clouds near the Galactic Center and above the Galactic plane, a detailed modeling of the cosmic-ray ionization contribution and, potentially, a more refined analysis of the gas density in clouds through dust extinction maps. Thus, ionization of molecular clouds emerges as one of the most powerful tools for probing sub-keV dark matter.
I Introduction
Despite compelling evidence for the existence of dark matter (DM) through its gravitational effects, the fundamental nature of DM remains largely unknown. If DM is indeed a new fundamental particle, it is crucial to explore its possible interactions beyond gravity; see Ref. Cirelli et al. (2024) for a comprehensive review. Numerous studies have placed stringent constraints on heavy fermionic (WIMP-like) DM with masses in the GeV range and above Aalbers et al. (2024); Bo et al. (2025); Aprile et al. (2025); Agnese et al. (2018); Agnes et al. (2023); Ajaj et al. (2019); Akerib et al. (2019). Significant efforts are currently underway to probe lighter DM candidates through both direct and indirect detection strategies Antel et al. (2023). These approaches aim to uncover signals from particles light enough that escape the traditional WIMP-focused searches. A wide array of theoretical models extending the Standard Model (SM) suggest the existence of light, weakly interacting particles that could account for some or all of the present-day DM density. These extensions often emerge in attempts to address unresolved issues in particle physics, such as the smallness of neutrino masses or the strong-CP problem. Among the most extensively studied candidates in this mass range there is the QCD axion Peccei (2008), a pseudoscalar particle originally proposed to solve the strong-CP problem, and more general axion-like particles (ALPs) Choi et al. (2021), which appear in many string-theoretic frameworks and span a broad range of masses and couplings Gendler et al. (2024). Significant effort is devoted to sterile neutrinos Dasgupta and Kopp (2021), which are singlets under the SM gauge group but may mix with active neutrinos and contribute to the DM and neutrino phenomenology, and dark photons Fabbrichesi et al. (2020), hypothetical gauge bosons associated with a hidden symmetry that can kinetically mix with the SM photon. The search for these light DM candidates employs a diverse set of experimental techniques. Direct detection experiments attempt to measure the scattering or absorption of such particles in terrestrial detectors Lanfranchi et al. (2021), while indirect detection methods aim to observe their decay or annihilation products in astrophysical environments Raffelt (1996); Caputo and Raffelt (2024); Jacobsen et al. (2023); Foster et al. (2022); Carenza et al. (2025).
In this work, we follow this latter path and propose a new method to probe the coupling of DM with masses at the keV scale and below. In particular, we show that the annihilation/decay of such low-mass DM particles into photons from ultraviolet (UV) to X-ray energies is able to produce significant ionization inside molecular clouds (MCs), reaching levels of ionization even above the ionization that cosmic rays (CRs) are expected to cause. Interactions between DM, or its annihilation/decay products, and the visible sector could ionize molecular hydrogen, , which is abundant within MCs. For example, DM decay into photons can lead to ionization when the DM mass satisfies eV, since the ionization energy of molecular hydrogen is approximately eV, producing high levels of ionization when these photons are absorbed inside the MC. Similarly, a coupling between DM and electrons can cause ionization through scattering processes. In dense molecular media, and, in particular, in dense MCs, UV photons are quickly absorbed, delivering all of their energy into the medium and ionizing the gas. Since DM would be nearly homogeneously distributed inside these clouds (extending over scale order of parsecs), its annihilation/decay would lead to an intense, homogeneous, ionization inside such systems. This potential DM-induced ionization contributes alongside standard ionization sources, such as stellar UV radiation, low-energy CRs, and X-rays emitted by stars within the MC Krolik and Kallman (1983); Norman and Silk (1980). It is possible to set strong constraints on the DM lifetime, annihilation rate, or on the coupling to photons when DM generates a ionization that exceeds the standard astrophysical sources.
Previous literature already investigated the heating of gas Bhoonah et al. (2018); Wadekar and Farrar (2021); Alvey et al. (2023) or the ionization Blanco et al. (2024); Balan et al. (2024) induced by the direct interaction of DM particles with the nuclei in gas systems, such as rich-gas dwarf galaxies, MCs or the early universe, through the cm line Evoli et al. (2014) for example. However, the effect of particle emissions from DM in the ionization of MCs has not been investigated before, making this the first effort to illustrate the potential of such clouds to indirectly probe light DM. As we show, on astrophysical scales, MCs act as large-volume detectors with ultra-low thresholds for ionization effects.
This paper is organized as follows: In section II we overview the expected main drivers of ionization in MCs and briefly outline the measurements of the ionization rate. Then, we discuss the targets that we are going to use in our study and identify optimal MCs to constraint DM. In Sec. III we detail our calculation of the ionization caused by photons produced from DM decay/annihilation. Then, we describe our results is Sec. IV and summarize our main findings in Sec. V, where we discuss possible ways of optimizing our current constraints.
II Ionization of dense molecular clouds
@@
While MCs with low column densities111The column density is defined as the integral of the density of a system over our line of sight, , N cm-2, called diffuse MCs, are expected to be highly affected by external UV radiation penetrating the cloud, above this column density threshold, corresponding to dense MCs Snow and McCall (2006), the UV radiation is strongly attenuated, and low-energy CRs are expected to be the dominant ionizing agent McKee (1989). In fact, CRs are expected to drive the chemical complexity observed in neutral gas systems and play a major role in the regulation of star formation. Modeling the contribution of CRs to the ionization of MCs is an intricate problem. Early studies estimated that the ionization rate per atom in clouds composed mainly of atomic hydrogen is within the range of a few Spitzer and Tomasko (1968); Indriolo and McCall (2012); Phan et al. (2023); Indriolo et al. (2015). This broad uncertainty reflects poor knowledge of the CR spectra at energies much below MeV. In particular, the ionization cross section for atomic hydrogen peaks at approximately eV for CR electrons and around keV for CR protons Padovani et al. (2009). In order to determine with more accuracy the expected ionization rates, a better understanding of the low-energy spectra of CR electrons, protons, and heavier nuclei is required. This is an extremely challenging task. The simplest approach consists of measuring the low-energy CR spectrum reaching Earth and extrapolating it to energies below the CR observations. Unfortunately, local measurements of the CR spectra at low energies are affected by solar modulation and the structure of the interplanetary magnetic field. The Voyager I and II spacecrafts have provided valuable data, measuring the electron spectrum down to a few MeV Cummings et al. (2016); Stone et al. (2019) beyond the heliopause, although these observations are still far from the energy range most relevant for ionization. An alternative to probing these low energies involves observations of the Galactic synchrotron background. This method is limited by significant uncertainties in modeling the interstellar magnetic field Padovani and Galli (2018); Orlando (2018); Ginzburg and Syrovatskii (1965) and it would hardly probe energies of tens of eV without involving a low-energy extrapolation of the CR spectrum. Thus, modeling the ionization of dense MCs is subject to significant uncertainties.
@@
Experimentally, measuring the ionization of MCs is of interest. A particularly effective method to infer the CR ionization rate involves measuring the intensities of emission lines from molecular tracers, whose abundances are closely tied to the free electron population. The choice of tracer strongly depends on the gas density and chemical composition. In diffuse regions of MCs, is a favorable tracer due to its simple and well-understood chemistry Indriolo and McCall (2013). In denser environments, such as MC cores and proto-stellar clusters, commonly used tracers include HCO+, DCO+, and CO Caselli et al. (1998). At present, measurements of the ionization rate in MCs seem challenging to explain by CR ionization. First, if CRs are the dominant ionization agent of MCs across the Galaxy, one expects that the ionization rate decreases for MCs with larger column densities. Although there are observations that support this trend, no clear correlation is observed Gabici et al. (2019); Gabici (2022). Second, the levels of ionization from CRs are expected to be lower than that measured in several clouds Indriolo et al. (2009); Phan et al. (2018). In this context, the high ionization rate observed in the Central Molecular Zone have been found to be incompatible with what the ionization that CRs can provide Ravikularaman et al. (2025), and MeV DM has been proposed as a potential explanation of this anomalous ionization De la Torre Luque et al. (2025). In addition, there are hints from gamma-ray observations suggesting that CRs do not actually penetrate dense MCs Yang et al. (2023).
As opposed to standard ionization mechanisms, UV photon produced by DM interactions would ionize directly within the cloud, providing a constant additional ionization to that provided by CRs. In this work, we explore the potential of dark, or dense, clouds, which are observed with temperatures near K, sizes from a below the parsec to tens of parsecs and masses up to M⊙ Bergin and Tafalla (2007), to probe the emission of UV/X-ray photons from sub-keV DM. In order to show how promising are such clouds as targets for DM searches, we will derive constraints from a local dark cloud as a pessimistic (and robust) case, because the modeling of the DM density does not need assuming any DM density distribution and will therefore be very robust. Then, we will evaluate constraints from a cloud closer to the Galactic Center (GC), where DM is expected to be concentrated, and we will assume a prototypical Navarro-Frenk-White (NFW) Navarro et al. (1996) distribution. This is the standard assumption made in most of the indirect DM searches, and allows for a fair estimation with previous constraints. As we will explain below, we follow a very conservative approach in both cases, and better results can be obtained through a more dedicated search of optimal clouds that is left for future work.
@@
We stress that L1551 is not the local cloud with the lowest ionization rate. For example, the Taurus molecular cloud (TMC-1) is a dark cloud in the constellations of Taurus and Auriga. For this cloud, Ref. Flower et al. (2007) found that provided the best overall agreement with the observations. Using this cloud would improve our limits by a factor on the DM annihilation/decay rate. However, the lack of a robust estimation of the uncertainties of this measurement makes us choose L1551 as reference case. Another interesting example is Barnard 68, which was found to be compatible with an ionization rate of Maret and Bergin (2007). Unfortunately, there is no precise estimation of the uncertainty of these measurements that allow us to use them reliably, but it must be noted that using these clouds could lead to up to an order of magnitude stronger DM constraints than our reference case. Lower values of ionization can be found in the local environment, but we leave a dedicated survey for a future work.
@@
Interestingly, even lower values of ionization have been found in infrared dark clouds (IRDC), which are very cold (T K) and dense clouds, and can be ideal targets for DM searches. Among known IRDCs, one that has been very well studied recently is G28.37+00.07, often referred to as the DRAGON cloud, which is located around kpc away from the Galactic center (in the so-called -kpc ring) Simon et al. (2006). This cloud has a size around L pc, a density of n cm-3 and a column density N cm-2. Ref. Entekhabi et al. (2022) studied the ionization rate and other gas properties across different sub-regions of this cloud, finding that the ionization rate in some sub-regions was exceptionally low. Indeed, the authors of this paper chose some of these sub-regions because they were expected to gather no star formation activity. Since the ionization from photons produced by DM is expected to be uniform throughout the cloud, due to negligible variations in DM density on scales of a few parsecs, constraints on DM-induced ionization can be derived from measurements in individual zones within the cloud. In particular the region P8 of that study was found to have ionization values between s-1 and s-1 in their fiducial model. We set an upper limit on the ionization of this cloud by taking twice the highest value found in this analysis, i.e. s-1. This will correspond with our fiducial constraint.
@@
While the modeling of DM ionization in local clouds would not require to assume any DM distribution (i.e. leading to really robust bounds), we select a cloud located quite close to the GC to demonstrate the potential of such dark clouds to probe sub-keV DM. The advantage is the higher DM density in this region, with the disadvantage of introducing uncertainties related to the DM distribution profile. For concreteness, we choose G1.4-1.8+87, with a size of around pc, and located around pc away from GC, above the Galactic plane Bhoonah et al. (2018). This cloud was found to be compatible with an ionization rate as low as Bhoonah et al. (2019), however, Ref. Farrar et al. (2020) casted serious doubts on the gas temperature used in this analysis. We will use this cloud as an example of the optimistic case of MC with low ionization rate located close to the Galactic Center. However, this should not be interpreted as a strict constraint, but rather as a forecast to illustrate the potential of similar clouds once more robust analyses are performed. To derive this limit, we choose a value of ionization of and will assume a prototypical NFW DM profile, which is the standard assumption in this kind of searches. We will call this an optimistic constraint, that will be shown in a different way than the fiducial and conservative cases, for clarity. As in the case of local MCs, more research is needed to identify more clouds close to the GC. IRDCs located close to the GC but above the Galactic plane are probably the best target to constraint the ionization caused by sub-keV DM, since the impact of CR ionization will be significantly reduced.
III Ionization of UV/X-ray photons from keV DM
@@
The ionization rate , for a MC located at x caused by photons with energy , can be calculated as Padovani et al. (2009); Phan et al. (2023)
@@
Here, eV is the ionization potential of H2, is the energy of the electron produced from the photoionization induced by an incoming photon with energy and eV is the average energy loss of a non-thermal electron per ionization of an H2 molecule in the interstellar medium Dalgarno et al. (1999). The integral in Eq. 1 extends from a minimum energy of to the maximum photon energy produced in the DM annihilation/decay, i.e. the mass of the DM particle, mχ, in the case of annihilating DM and mχ/2 in the case of decaying DM. Here, we are neglecting the Compton ionization process, which is important only above tens of keV Yan et al. (2001). This simplification is motivated by the fact that at such energies, our limits are not competitive with previous constraints.
@@
In Fig. 2 we show the quantity J for several DM masses in a local MC like L1551. As expected, the lower the DM mass, the higher is J and, thus, the ionization rate. This behavior is due to the higher DM number density and the higher photoionization cross sections, which peaks around eV Yan et al. (2001). This can be also seen from Fig. 2, where we show the H2 ionization rate caused by annihilating (dot-dashed line) and decaying (solid lines) DM in a local MC. The purple dashed line shows the level of ionization observed for L1551. In this work we will constrain the annihilation rate into photons defined as , where is the total DM annihilation rate and is the branching ratio for photons in the final state. Similarly, we will always refer to DM decay into photons.
@@
For the local cloud L1551, we set the DM density to be the local DM density, which is GeV/cm3 Benito et al. (2021) at the Earth position kpc from the Galactic center Evans et al. (2019). We stress that this makes our estimation independent on the DM density profile, which is one of the main uncertainties in DM indirect searches. In fact, given that the considered cloud is located in the Solar System neighborhood, this assumption leads to a quite conservative result. For the case of the DRAGON cloud, which corresponds to our fiducial constraint, we assume the NFW profile Navarro et al. (1996), as it is the most usual assumption used for DM searches. Considering constraints on the DM lifetime , using the NFW profile leads to a factor stronger bound than assuming the local DM density at the cloud position, while for annihilation () the difference is close to a factor . We stress that cuspier DM profiles, like the ones favored by the GC excess Ackermann et al. (2017), are commonly used as optimistic scenarios, but we will not consider them. We also adopt a NFW profile for our optimistic forecast with G1.4-1.8+87. We also note that the local DM density value that we have taken is in the lower bound given in Ref. Benito et al. (2021), where the estimated local DM density is GeV/cm3 at 2 level.
IV Results
@@
In Fig. 3, we report our model-independent constraints on DM two-photon decay (left panel) and annihilation (right panel). In this energy range, the strongest constraints are from CMB spectral distortions Bolliet et al. (2021) and heating of gas-reached dwarf galaxies (particularly, the Leo T dwarf) Wadekar and Wang (2022). The latter constraints not only requires the adoption of a model for the DM distribution, gas temperatures and the density distributions of molecular and atomic gas phases as well as dust grains (which are uncertainties that can similarly affect the measurement of the ionization rate of the DRAGON cloud), but it also requires modeling the background heating from CRs and UV/X-ray. Thus, different assumptions imply different uncertainties that are involved in these constraints. Here, we show that even in the most pessimistic case of a local MC such as L1551, with a low DM density, the constraints at tens of eV are competitive with the existent bounds Wadekar and Wang (2022); Bolliet et al. (2021); Audren et al. (2014); Poulin et al. (2016). Notably, our fiducial bound (the DRAGON cloud), becomes stronger than CMB constraints in the mass range from eV to almost keV, and it is also stronger than the LeoT gas heating constraints up to a few hundreds eV. As commented above, the assumption of a NFW profile for the DRAGON cloud makes our constraint only a factor of stronger than assuming only the local DM density at the cloud position, and around a factor of for the annihilation rate constraint. Taking a Burkert profile Burkert (1995), which is often taken as the most pessimistic scenario in Galactic DM searches, these limits are expected to be reduced by a factor for decay and for annihilation. We remind the reader, that even cuspier profiles than NFW are equally likely, and those would result in an improvement of this limit. Notably, for the optimistic forecast with G1.4-1.8+87 we get a constraint that can only be improved by direct X-ray observations, such as those from XMM-Newton Gewering-Peine et al. (2017); Foster et al. (2021) above the keV. We tested another interesting candidate, the cloud G357.8-4.7-55, which is around kpc Bhoonah et al. (2018) far from the Galactic Center and is also compatible with low ionization rates. Assuming a NFW, we get a constraint that is more than one order of magnitude better than the one from L1551 for the lifetime constraint.
@@
We have also estimated what could be the effect of adding the ionization from CRs in L1551. Since CRs are expected to be the ionization agents in the interstellar medium, their effect will be dominant. In fact, previous predictions expect the ionization of L1551 to be totally explained by CRs Phan et al. (2023). To remain conservative in our estimation, we adopt the CR ionization calculated in Ref. Phan et al. (2023), pointing towards an ionization rate between s-1 and s-1 in the diffusive regime. This will be the contribution to the ionization of L1551 explained by standard physics. We adopt the most conservative value of this estimation, s-1 in our calculation. We find that adding this contribution leads to around a factor of better constraints below keV for both decay and annihilation. This rough approach illustrates the improvement expected in the derived limits when accounting for the background ionization of CRs. A similar estimation could be made for clouds that are located farther away from the Solar System, but this would require a precise modeling of the distribution of low-energy CRs, and the uncertainties would be very important.

### v2

@@ Abstract
The ionization of dense molecular clouds offers a powerful and robust probe of dark matter models that produce UV or X-ray photons via annihilation or decay. Using this novel observable, we derive competitive constraints on dark matter in the mass range from 30 eV to 1 keV. In particular, we set the strongest existing limits on axion-like dark matter in the 30–100 eV range. We also present sensitivity projections demonstrating the significant potential of this approach. These constraints can be substantially improved with targeted observations of molecular clouds near the Galactic Center and above the Galactic plane as well as with a more accurate modeling of cosmic-ray ionization and refined gas density estimates from dust extinction maps. Our results establish molecular cloud ionization as one of the most sensitive tools for probing sub-keV dark matter.
Introduction — Uncovering the fundamental nature of dark matter (DM) remains one of the most pressing challenges in physics. Therefore, it is essential to investigate possible non-gravitational interactions of DM with the visible sector Cirelli et al. (2024). Light and weakly coupled particle models are attracting significant attention Lanfranchi et al. (2021); Antel and others (2023) because of their potential in addressing other open problems, such as the strong-CP problem. The astrophysical signatures of light DM models are unique, motivating a wide range of indirect detection strategies Raffelt (1996); Caputo and Raffelt (2024); Jacobsen et al. (2023); Foster et al. (2022); Carenza et al. (2025b); Nguyen et al. (2025).
In this work we identify the ionization of dense molecular clouds (MCs) as a new and highly sensitive astrophysical probe of photons injected by light DM annihilation or decay. The photons injected by sub-keV DM have energies ranging from ultraviolet (UV) to X-ray, and they are able to produce significant ionization of molecular hydrogen, in MCs. DM-induced ionization might even exceed the ionization that cosmic rays (CRs) are expected to cause in a standard scenario Krolik and Kallman (1983); Norman and Silk (1980). Moreover, DM is a ionization source uniformly distributed across the MC, injecting ionizing radiation directly inside the dense MC. When photons are injected by DM in a dense MC, they are quickly absorbed and efficiently ionize the medium.
In the sub-keV mass range, there are many well-motivated extensions of the Standard Model (SM) predicting DM particles able, for example, to generate photons in their decay. Axion-like particles (ALPs), including the QCD axion Peccei (2008) as a special case, are generically predicted to couple weakly to photons and possibly contribute to DM Choi et al. (2021); Gendler et al. (2024). The simplest and most motivated QCD axion models Peccei and Quinn (1977); Weinberg (1978); Wilczek (1978) can easily produce cold DM through the misalignment mechanism Preskill et al. (1983); Abbott and Sikivie (1983); Dine and Fischler (1983) or topological defect decay Harari and Sikivie (1987); Hagmann and Sikivie (1991); Buschmann et al. (2022); Kawasaki et al. (2015). The robust stellar bounds Ayala et al. (2014); Carenza et al. (2020); Lucente et al. (2022) usually preclude the simplest realizations of a QCD action in the eV-keV mass range. However, assuming the existence of degenerate SM copies, with the axion being the common portal in all the replicas Di Luzio et al. (2021); Giannotti (2005); Di Luzio et al. (2020), increases the QCD axion mass up to the eV-keV scale. Another possibility is that the axion couples to a copy of the QCD sector, which introduces a dynamical energy scale around GeV Rubakov (1997), increasing the axion mass at fixed photon coupling. See also Ref. Gavela et al. (2024) for models with many axions and Ref. de Giorgi and Ramos (2025) for models with extra dimensions.
The emergence of ALPs in the eV-keV mass range might be unrelated to solving the strong CP problem, and for this reason we refer to them as ALPs and not QCD axions. As a matter of fact, ALPs appear as pseudo-Nambu-Goldstone bosons of a broken global symmetry. Thus, their mass and couplings will be unrelated. The interactions between ALPs and photons lead to the production of ALPs from the thermal plasma in the early Universe. In the ALP mass range between tens of eV and keV there are stringent constraints requiring the ALP-photon coupling to be . For such a small coupling, ALPs will never reach thermal equilibrium with the visible sector, leaving the freeze-in scenario as a possibility. Non-thermal production mechanisms, such as misalignment and topological defect decay O’Hare et al. (2022); Gelmini et al. (2021), are a more appealing possibility. The misalignment mechanism, is unlikely to allow for ALP DM in the eV-keV range unless some fine-tuning on the initial conditions is involved. By contrast, depending on the details of the topological defect annihilation, ALPs in the considered mass range might compose the entirety of DM, that leave observable gravitational wave imprints Gelmini et al. (2021). Several QCD axion and ALP models predict viable DM in the eV–keV range, motivating searches in this regime.
In this Letter, we prove that the MC ionization bound can impose nearly leading constraints on axion DM from 30 eV to 200 eV, as illustrated in Fig. 1, with extremely good perspectives for future improvements. In Fig. 1, we compare existent limits with the bounds from different, well characterized, MCs (L1551 and the DRAGON cloud), and we show an optimistic forecast for a MC near the GC, G1.4-1.8+87, as described below. This exemplary DM model proves the great potential of using MCs as ultrasensitive natural detectors of DM-induced ionization.
@@
Ionization of dense molecular clouds — While MCs with low column densities111The column density is defined as the integral of the density of a system over our line of sight, , N cm-2, called diffuse MCs, are expected to be highly affected by external UV radiation penetrating the cloud, above this column density, corresponding to dense MCs Snow and McCall (2006), the UV radiation is strongly attenuated, and low-energy CRs are expected to be the dominant ionizing agent McKee (1989). In fact, CRs are expected to drive the chemical complexity observed in neutral gas systems and play a major role in the regulation of star formation. Modeling the contribution of CRs to the ionization of MCs is an intricate problem. Early studies estimated that the ionization rate per atom in clouds composed mainly of atomic hydrogen is within the range of a few Spitzer and Tomasko (1968); Indriolo and McCall (2012); Phan et al. (2023); Indriolo et al. (2015). This broad uncertainty reflects poor knowledge of the CR spectra at energies much below MeV.
@@
Experimentally, measuring the ionization of MCs is of interest. A particularly effective method to infer the CR ionization rate involves measuring the intensities of emission lines from molecular tracers, whose abundances are closely tied to the free electron population. The choice of tracer strongly depends on the gas density and chemical composition. In diffuse regions of MCs, is a favorable tracer due to its simple and well-understood chemistry Indriolo and McCall (2013). In denser environments, such as MC cores and proto-stellar clusters, commonly used tracers include HCO+, DCO+, and CO Caselli et al. (1998). At present, measurements of the ionization rate in MCs seem challenging to explain by CR ionization. First, if CRs are the dominant ionization agent of MCs across the Galaxy, one expects that the ionization rate decreases for MCs with larger column densities. Although there are observations that support this trend, no clear correlation is observed Gabici et al. (2019); Gabici (2022). Second, the levels of ionization from CRs are expected to be lower than that measured in several clouds Indriolo et al. (2009); Phan et al. (2018). In this context, the high ionization rate observed in the Central Molecular Zone have been found to be incompatible with what the ionization that CRs can provide Ravikularaman et al. (2025), and MeV DM has been proposed as a potential explanation of this anomalous ionization De la Torre Luque et al. (2025). In addition, there are hints from gamma-ray observations suggesting that CRs do not actually penetrate dense MCs Yang et al. (2023). As opposed to standard ionization mechanisms, UV photon produced by DM interactions would ionize directly within the cloud, providing a constant additional ionization to that provided by CRs. In this work, we explore the potential of dark, or dense, clouds, which are observed with temperatures of tens of Kelvin, sizes from a below the parsec to tens of parsecs and masses up to M⊙ Bergin and Tafalla (2007), to probe the emission of UV/X-ray photons from sub-keV DM. In order to show how promising are such clouds as targets for DM searches, we will derive constraints from a local dark cloud as the most pessimistic case, because the modeling of the DM density does not need assuming any DM density distribution and will therefore be very robust. Then, we will evaluate constraints from a cloud closer to the Galactic Center (GC), where DM is expected to be concentrated, and we will assume a prototypical Navarro-Frenk-White (NFW) Navarro et al. (1996) distribution.
@@
We stress that L1551 is not the local cloud with the lowest ionization rate. For example, the Taurus molecular cloud (TMC-1) is a dark cloud in the constellations of Taurus and Auriga. For this cloud, Ref. Flower et al. (2007) found that provided the best overall agreement with the observations. Using this cloud would improve our limits by a factor on the DM annihilation/decay rate. However, the lack of a robust estimation of the uncertainties of this measurement makes us choose L1551 as reference case. Another interesting example is Barnard 68, which was found to be compatible with an ionization rate of Maret and Bergin (2007). Unfortunately, there is no precise estimation of the uncertainty of these measurements that allow us to use them reliably, but it must be noted that using these clouds would lead to at least a factor of a few stronger DM constraints than L1551. Lower values of ionization can be found in the local environment, but we leave a dedicated survey for a future work.
@@
Interestingly, similar or even lower values of ionization have been found in infrared dark clouds (IRDCs), which are very cold (T K) and dense clouds, and can be ideal targets for DM searches. Among known IRDCs, one that has been very well studied recently is G28.37+00.07, often referred to as the DRAGON cloud, which is located around kpc away from the GC (in the so-called -kpc ring) Simon et al. (2006). This cloud has a size around L pc, a density of n cm-3 and a column density N cm-2. Ref. Entekhabi et al. (2022) studied the ionization rate and other gas properties across different sub-regions of this cloud, finding that the ionization rate in some sub-regions of the cloud was exceptionally low. Indeed, the authors of this paper chose some of these sub-regions because they were expected to gather no star formation activity. Since the ionization from photons produced by DM is expected to be uniform throughout the cloud, due to negligible variations in DM density on scales of a few parsecs, constraints on DM-induced ionization can be derived from measurements in individual zones within the cloud. In particular the region P8 of that study was found to have ionization values between s-1 and s-1 in their fiducial model. We set an upper limit on the ionization of this cloud by taking twice the highest value found in this analysis, i.e. s-1, which corresponds to our fiducial constraint. The authors of the same paper also performed an additional analysis that included more ionized species, noting that these species have more uncertain astrochemistry and are less directly linked to the induced ionization rates, which may introduce additional uncertainty into their analysis. This analysis yielded ionization rates of and . Accordingly, we present the results for the DRAGON cloud as a band, adopting constraints in the range to The P8 region has a size of pc, with a gas density around cm-3.
@@
While the modeling of DM ionization in local clouds would not require to assume any DM distribution, we also select a cloud located quite close to the GC to demonstrate the potential of such dark clouds to probe sub-keV DM. The advantage is the higher DM density in this region, with the disadvantage of introducing uncertainties related to the DM distribution profile. For concreteness, we choose G1.4-1.8+87, with a size of around pc, and located around pc away from GC, above the Galactic plane Bhoonah et al. (2018). This cloud was found to be compatible with an ionization rate as low as Bhoonah et al. (2019), however, Ref. Farrar et al. (2020) casted serious doubts on the gas temperature used in this analysis. We will use this cloud as an example of the optimistic case of MC with low ionization rate located close to the GC. However, this should not be interpreted as a strict constraint, but rather as a forecast to illustrate the potential of similar clouds once more robust analyses are performed. To derive this limit, we choose a value of ionization of and will assume a prototypical NFW DM profile, which is the standard assumption in this kind of searches. We will call this an optimistic constraint, that will be shown in a different way than the fiducial and conservative cases, for clarity. As in the case of local MCs, more research is needed to identify more clouds close to the GC.
@@
Ionization of UV/X-ray photons from keV DM — The ionization rate , for a MC located at x caused by photons with energy , can be calculated as Padovani et al. (2009); Phan et al. (2023)
@@
Here, eV is the ionization potential of H2, is the energy of the electron produced from the photoionization induced by an incoming photon with energy and eV is the average energy loss of a non-thermal electron per ionization of an H2 molecule in the interstellar medium Dalgarno et al. (1999). Precisely, the values of W we taken from Fig. 6 of Ref. Dalgarno et al. (1999). The integral in Eq. 1 extends from a minimum energy of to the maximum photon energy produced in the DM annihilation/decay, i.e. the mass of the DM particle, mχ, in the case of annihilating DM and mχ/2 in the case of decaying DM. Here, we are neglecting the Compton ionization process, which is important only above tens of keV Yan et al. (2001). This simplification is motivated by the fact that at such energies, our limits are not competitive with previous constraints.
@@
For the local cloud L1551, we set the DM density to be the local DM density, which is GeV/cm3 Benito et al. (2021) at the Earth position kpc from the GC Evans et al. (2019). We stress that this makes our estimation independent on the DM density profile, which is one of the main uncertainties in DM indirect searches. For the case of the DRAGON cloud, which corresponds to our fiducial constraint, we assume the NFW profile Navarro et al. (1996), as it is the most usual assumption used for DM searches. We also adopt a NFW profile for our optimistic forecast with G1.4-1.8+87. We note that the local DM density value that we have taken is in the lower bound given in Ref. Benito et al. (2021), where the estimated local DM density is GeV/cm3 at 2 level.
The previous estimate assumes complete absorption of the UV photons emitted by DM within the cloud; however, this assumption does not always hold. We thus compute the opacity of the cloud () to photons emitted by DM. The resulting absorption probability, , is shown in Fig. S2 of the Supp. Material as a function of photon energy. In our calculations, specifically in Eq. 1, we multiply the integrand by this absorption probability to account for photons that escape the cloud, which suppresses our constraints towards high photon energies.
@@
Results — In Fig. 2, we report our model-independent constraints on DM two-photon decay (left panel) and annihilation (right panel). In this energy range, the strongest constraints are from CMB spectral distortions Bolliet et al. (2021b) and heating of gas-reached dwarf galaxies (particularly, the Leo T dwarf) Wadekar and Wang (2022). Here, we show that even in the most pessimistic case of a local MC such as L1551, with a low DM density, the constraints at tens of eV are competitive with the existent bounds Wadekar and Wang (2022); Bolliet et al. (2021b); Audren et al. (2014); Poulin et al. (2016). Notably, our fiducial bound (the DRAGON cloud), becomes stronger than CMB constraints in the mass range from eV to several hundreds eV, and it is also stronger than the LeoT gas heating constraints up to a hundred eV. As commented above, the assumption of a NFW profile for the DRAGON cloud makes our constraint only a factor of stronger than assuming only the local DM density at the cloud position, and around a factor of for the annihilation rate constraint. Taking a Burkert profile Burkert (1995), which is often taken as the most pessimistic scenario in Galactic DM searches, these limits are expected to be reduced by a factor for decay and for annihilation (see Fig. S2 in section S.I. of the Supp. Material, where the parameters characterizing the DM density distribution used are given). We remind the reader, that even cuspier profiles than NFW, like the ones favored by the GC excess Ackermann and others (2017) are equally likely, and those would result in an improvement of this limit. On top of this, we note that the presence of DM minihalos or clumps within these clouds will substantially improve the predicted ionization rates, therefore leading to much stronger constraints. We remain conservative by only considering the subjacent DM distribution. Notably, for the optimistic case with G1.4-1.8+87 we get a constraint that is far stronger than previous observations below a hundred eV above the keV. We tested another interesting candidate, the cloud G357.8-4.7-55, which is around kpc Bhoonah et al. (2018) far from the GC and is also compatible with low ionization rates. Assuming NFW profile, we get a lifetime constraint one order of magnitude stronger than the L1551 MC for decaying DM.
@@
We have also estimated what could be the effect of adding the ionization from CRs in L1551. Since CRs are expected to be the ionization agents in the interstellar medium, their effect will be dominant. In fact, previous predictions expect the ionization of L1551 to be totally explained by CRs Phan et al. (2023). We adopt the CR ionization calculated in Ref. Phan et al. (2023), pointing towards an ionization rate between s-1 and s-1 in the diffusive regime. We adopt the most conservative value of this estimation, s-1 in our calculation, and find that adding this contribution leads to around a factor of better constraints below keV for both decay and annihilation. This rough approach illustrates the improvement expected in the derived limits when accounting for the background ionization of CRs. A similar estimation could be made for clouds that are located farther away from the Solar System, but this would require a precise modeling of the distribution of low-energy CRs, and the uncertainties would be very important.